# Beyond Trial-and-Error: Predicting User Abandonment After a Moderation Intervention

### v2

@@ X Conclusions
We proposed and tackled —for the first time— the task of predicting the effects of a moderation intervention. Specifically, we detected those users who abandoned Reddit following the ban of a large number of communities on the platform. To solve this task we investigated the behavior of 16,540 users by leveraging 16M comments they posted in the banned communities and 13.8M comments they posted in non-banned ones. Starting from this extensive dataset, we extracted 142 features conveying information about the activity, toxicity, relations, and writing style of the analyzed users. Our results are promising, albeit preliminary as one would expect from a new task, with the best model achieving micro F1 $=0.800$ and macro F1 $=0.676$. Our model proved to be sufficiently generalizable when applied to users from unseen communities. Furthermore, we found that activity features are the most informative for this task, followed by relational and toxicity features. Conversely, writing style features seem to provide limited information. Given the novelty of the task, promising directions for future work are multifold. Among them are the inclusion of additional features that could provide relevant information on user behavior, and particularly, on their possible reactions to a moderation intervention, such as socio-psychological features. Additionally, one could cast the problem of predicting the effects of a moderation intervention as a regression task, rather than as a classification task as done in the present work. To this regard, being able to precisely estimate the behavioral changes of some users would provide even more fine-grained information to platform administrators for planning their moderation actions.
@@ Acknowledgments
This work is partially supported by the European Union, Next Generation EU, within the PRIN 2022 framework project PIANO (Personalized Interventions Against Online Toxicity); by the PNRR-M4C2 (PE00000013) ”FAIR-Future Artificial Intelligence Research” - Spoke 1 ”Human-centered AI”, funded under the NextGeneration EU program; and by the Italian Ministry of Education and Research (MUR) in the framework of the FoReLab project (Departments of Excellence).

### v3

@@ 10 Conclusions and Future Work
We proposed and tackled —for the first time— the task of predicting the effects of a moderation intervention. Specifically, we detected those users who abandoned Reddit following the ban of a large number of communities on the platform. To solve this task, we investigated the behavior of 16,540 users by leveraging 16 million comments they posted in the banned communities and 13.8 million comments they posted in non-banned ones. Starting from this extensive dataset, we extracted 142 features conveying information about the activity, toxicity, relations, and writing style of the analyzed users. Our results are promising, albeit preliminary as one would expect from a new task, with the best model achieving micro F1 $=0.914$. Our model proved to be sufficiently generalizable when applied to users from unseen communities. Furthermore, we found that activity features are the most informative for this task, followed by relational and toxicity features. Conversely, writing style features seem to provide limited information.
Given the novelty of the task, promising directions for future work are multifold. Among them are the inclusion of additional features that could provide relevant information on user behavior, and particularly, on their possible reactions to a moderation intervention, such as socio-psychological features. Other promising features are those that allow tracking linguistic and semantic shifts over time, including features based on sentiment, topics, and linguistic diversity. Future research could also focus on improving generalizability through the adoption of new datasets, enhancing scalability, enabling real-time applications, and ensuring multi-platform compatibility. Additionally, studying longitudinal changes in user behavior and their effects on abandonment predictions over time could provide deeper insights into the long-term impact of moderation interventions. Other favorable avenues of experimentation involve the adoption of more powerful machine learning methods, such as deep learning-based techniques (e.g., LSTMs or transformers). Finally, the task of predicting the effects of a moderation intervention could be cast as a multi-class, regression, or quantification problem, rather than as a binary classification as done in the present work. In this regard, being able to precisely estimate the extent of behavioral change exhibited by affected users would provide even more fine-grained information to platform administrators for planning their moderation actions.
@@ Acknowledgments
This work is partially supported by the European Union, Next Generation EU, within the PRIN 2022 framework project PIANO (Personalized Interventions Against Online Toxicity); by the PNRR-M4C2 (PE00000013) "FAIR-Future Artificial Intelligence Research" - Spoke 1 "Human-centered AI", funded under the NextGeneration EU program; and by the Italian Ministry of Education and Research (MUR) in the framework of the FoReLab project (Departments of Excellence).
[…]